OpenStudy (anonymous):

find the exact value of the expression Cos(sin^-1 square root(3) divided by 2)

OpenStudy (anonymous):

does it help to know that \(\cos^{-1}(\frac{\sqrt{3}}{2})=\frac{\pi}{3}\) ?

OpenStudy (anonymous):

im not sure how that helps..?

OpenStudy (anonymous):

\[\cos(\sin^{-1} \sqrt{3}\div2)\]